# Sub-millisecond electric field sensing with an individual rare-earth doped ferroelectric nanocrystal

### v1

@@ Abstract
Understanding the dynamics of electrical signals within neuronal assemblies is crucial to unraveling complex brain function. Despite recent advances in employing optically active nanostructures in transmembrane potential sensing, there remains room for improvement in terms of response time and sensitivity. Here, we report the development of such a nanosensor capable of detecting electric fields with a submillisecond response time at the single particle level. We achieve this by using ferroelectric nanocrystals doped with rare earth ions producing upconversion (UC). When such a nanocrystal experiences a variation of surrounding electric potential, its surface charge density changes, inducing electric polarization modifications that vary, via converse piezoelectric effect, the crystal field around the ions. The latter variation is finally converted into UC spectral changes, enabling optical detection of electric potential. To develop such a sensor, we synthesized erbium and ytterbium-doped barium titanate crystals of size nm. We observed distinct changes in the UC spectrum when individual nanocrystals were subjected to an external field via a conductive AFM tip, with a response time of 100 s. Furthermore, our sensor exhibits a sensitivity to electric fields of only 4.8 kV/cm/, making possible the time-resolved detection of a neuron action potential.
@@ 1 Introduction
The generation and propagation of electrical signals in neurons and across neuronal assemblies, which occurs at the nanometer and millisecond scales, remains poorly understood, despite the development of several theoretical models. For instance, the theory of core conductor cable and the Goldman-Hodgkin-Huxley-Katz models of ion motion and voltage propagation have provided theoretical foundations for neuronal physiology 1, 2. These theories account for the action potential, the few millisecond lasting voltage variation of 100 mV across the nm thick membrane (corresponding to 200 kV/cm peak voltage), which supports neuron-to-neuron communication. However, these models neglect electrodiffusion, that is, concentration changes associated with ionic currents. This assumption is true at millimeter scale like the squid giant axon but fails to describe accurately the processes taking place at the sub-micrometer size structure of synapses 3, 4. The Poisson-Nernst-Planck theory reveals that the geometry of synaptic compartments influences the distribution of local electric field generated during synaptic transmission, which subsequently changes the concentration of charged neurotransmitters undergoing electrodiffusion 5 and therefore has a functional impact. As the shape of the synapse changes during the execution of cognitive tasks 6, being able to measure the electric field on this scale would not only be an observable of such key plasticity but could provide an experimental test of the theory. There is thus a clear need to develop non-toxic, local methods to sense electric fields in biological media in order to refine current models of neuronal activity.
The detection of electric fields at the nanometer scale has initially been addressed using ultrahigh resolution nanopipette electrophysiology 7. Although very powerful, this technique can only study a single synapse at a time. On the contrary, voltage indicators (VI) were designed to stain the entire neuronal membrane and allow fluorescence recording of transmembrane voltage dynamics 8 throughout the cell with optical diffraction resolution. Enormous progress has been made in the development of brighter and faster VI, and in particular thanks to the development of hybrid chemigenetic systems that allows to combine optimally chemical compounds with genetically encoded proteins 9. However, subdiffraction imaging of a large number of synapses at high temporal resolution has not yet been reported. The development of a technology offering the necessary high spatiotemporal resolution (typically 1 ms, 200 nm) to reliably interrogate electrophysiological dynamics at multiple neuronal nanodomains would provide invaluable insights into open questions in neuronal electrophysiology.
@@
However, to maintain diffraction-limited imaging, an electric field must be sensed by a single or very few nanoparticles only. There are a few examples of such systems that are based on electric-field modulation of charge or Förster resonant energy transfer (FRET) between the nanoparticle and another entity. For example, Nag et al. used a QD-fullerene (C60) conjugate, with the C60 embedded in the cell membrane 12. At resting membrane potential the photoexcited QDs relax to the ground state partially by electron transfer to the C60. This transfer increases upon cell membrane depolarization, leading to a PL intensity relative quenching of a few % that provided a transmembrane potential variation readout both in cultured cell and in the cortex of live mouse. While chemical or electrical stimulation leads to QD PL modulation, the fastest response time was about 100 ms, two order of magnitude longer than the time response necessary to reproduce an action potential (AP) with fidelity. Another conjugate introduced by Liu et al. consists of an up-converting nanoparticle (UCNP) to which voltage sensing dyes were coupled, that serve on one hand to anchor the UCNP into cell membrane and are also FRET acceptors 13. The authors of this study report a relative change in UCNP luminescence intensity in the 0.4-6.1% range in cultured cells, depending on the stimulation conditions, and a time response of about 50 ms, still much longer than the duration of the AP. Hence, the spatial ( nm) and time (submillisecond) resolution required for membrane potential sensing remains to be addressed.
@@
Here, we report the development of a nm-sized standalone optically active nanocrystals (NCs) capable of sensing, at the level of a single particle, a field of kV/cm smaller than the peak amplitude field of an action potential), with a millisecond response time. We based our sensing principle on a specific UCNP made of a ferroelectric nanocrystal (NC) host doped with rare earth ions. When the NC is exposed to an external electric field, its lattice undergoes deformations due to a converse piezoelectric effect, which leads to spectral changes of the rare-earth ion up-conversion emission. The effectiveness of this principle was established by Hao et al. who demonstrated the modulation, by a varying electric field of 125 kV/cm amplitude, of the upconverted (UC) light from an ytterbium (Yb3+) and erbium (Er3+)-doped epitaxial barium titanate (BTO) thin film 14, at room temperature (in BTO ferroelectric phase). The authors demonstrated an UC modulation at a frequency of 0.1 Hz, far below the maximum frequency of kHz limited by the UC process response time of s in erbium-doped BTO 15. Thus, here we develop BTO:Yb3+,Er3+ individual nanocrystals of size nm, which we exposed to a fast varying electric field with the conductive tip of an atomic force microscope (AFM), coupled to a fluorescence microscope. We demonstrate that 29% of the nanocrystals exhibit a variation in UC intensity with a response time of s, limited by the duration of the UC optical cycle. Furthermore, BTO nanocrystals are ideal to develop sensors that could be rapidly tested in biological contexts due to their pre-existing use in drug delivery or tissue engineering 16, but also for their intrinsic optical two-photon excitation response17 (resulting from the crystal second-order nonlinear response) in cultured cells 18, 19, 20 and larvae of small organisms 21, 22, without or with grafted functional groups 22, 19, 20.
@@ 2.1 Synthesis and characterization of erbium and ytterbium-doped BTO nanocrystals
Erbium and ytterbium-doped BTO nanocrystals (nanoBTO:Yb,Er) were synthesized by a combination of coprecipitation and hydrothermal methods. Ytterbium was used to sensitize erbium up-conversion. By adapting the Ba/Ti stoichiometry, we aimed at incorporating Er3+ and Yb3+ ions in A and B sites of the ABO3 barium titanate perovskite respectively, because the ytterbium ions have the closest radius to the one of titanium (B site) in the six coordinated state 23. We targeted the stoichiometry , ensuring charge compensation, as described in Materials and Methods. The choice of an erbium concentration of 1% was made based on Zhang et al.24, who showed that higher concentrations enhance cross-relaxation, leading to lower UC.
@@
One may wonder where Er3+ and Yb3+ ions insert themselves in the BTO nanocrystal. Previous works 29, 30 indicate that Yb3+ ions predominantly substitute Ti4+ ions; for charge compensation reasons, it is thus highly likely that Er3+ substitutes Ba2+. To further investigate effective substitution sites, other studies, such as UC lifetime measurements 15 would be necessary but are beyond the scope of this article.
The characterization studies validated that the synthesized BTO nanocrystals are doped with ytterbium and erbium, and keep a tetragonal structure of undoped nanocrystals. Hence, nanoBTO:Yb,Er are very likely to maintain the ferroelectric and piezoelectric properties on which we based the electric field sensing mechanism. We then investigated the upconversion photoluminescence (PL) of individual nanoBTO:Yb,Er, and the modulation of its intensity by an external electric field.
@@ 2.2 Up-conversion of single nanoBTO:Yb,Er and its modification by an electric potential
Note that 4f-4f electric dipole transitions of lanthanide ions are expected to be forbidden by quantum mechanics selection rules in a perfectly centrosymmetric crystalline environment, for which the odd orders of the crystal field vanish. The mere detection of erbium UC, even in absence of an external electric field, is the signature of lower local symmetry around erbium ions. In fact, the splitting of the Miller indexes (002) and (200) Bragg peaks near in the inset of Fig. 1a already indicates the presence of the polar (hence non-centrosymmetric) phase of BTO, enabling the 4f-4f electric dipole transitions of lanthanide ions. Note also that Zou et al. reported erbium UC even in the high-temperature centrosymmetric phase of BTO 31, which indicates that the local environment felt by the RE ions are likely even less symmetric than the symmetry of the overall matrix, which is consistent with the eight-site model developed in BTO 32 in which local symmetry breakings exist at the unit cell level but vanish on a macroscopic scale.
@@
We then investigated the changes of UC spectrum of a single nanoBTO:Yb,Er (nanoBTO:Yb,Er #1) when we apply a static voltage of V across the particle. For the average particle size of 158 nm, this voltage corresponds to an electric field of kV/cm, more than twice smaller than the electric field breakdown of 1.5-3 MV/cm reported for thin BaTiO3 in the literature 33. Figure 4a shows UC spectral changes of nanoBTO:Yb,Er #1 under static voltage exposure of +10 V, manifested by an overall increase in the intensity of all peaks. This variation was reversible when the voltage was returned to 0 V, reverting to its initial value, and no change was detected with the opposite polarity of -10 V. For another NC (nanoBTO:Yb,Er #2), we also detected a reversible UC emission increase under an opposite applied voltage of -10 V (Figure 4b), also with no change for the opposite polarity (of +10 V). The intensities of the green and red UC emission bands both increase with the applied voltage, regardless of its polarity. This behavior differs from the one reported in BTO thin layer by Hao et al.14, where they observed an increase in the intensity of the green bands but none of the red one. Hao et al. proposed an interpretation based on the transition probability predicted by Judd–Ofelt theory, where the dominant contribution for the green emission transitions involves the intensity parameter that increases when the symmetry of the site occupied by Er3+ is reduced, which is what happens when the crystal is exposed to the external electric field. In contrary, the red emission transition does not involve this symmetry sensitive term and therefore should be almost insensitive to the electric field. Our observation of a variation of intensity of the red emission transition under an external electric field indicates that in our sample, there is a more complex coupling of the red transition to the site symmetry changes. We hypothesize that the change of intensity of the red line could be due to an increase of population of the 4F9/2 upper level of the red transition resulting from a larger cross-relaxation effect. The latter depends on the distance between ions, which indeed is modulated by the lattice deformation resulting from the converse piezoelectric effect, under the applied electric field. Indeed we evidenced that in our nanoBTO:Yb,Er sample the red transition is governed by the cross-relaxation, because we observed that its intensity evolves linearly with the excitation laser intensity (see Figure S3), consistently with two red photons emitted for two infrared excitation photons absorbed, with one of them experiencing a cross relaxation (see Figure 3d).
@@
As for NC like nanoBTO:Yb,Er #1 and #2 the intensity of both the green and red bands increases similarly under applied electric field, we considered the total UC intensity integrated over the whole spectrum. Figure 4c displays this total UC intensity change under a voltage square modulation for nanoBTO:Yb,Er #1, if whole the UC light is sent to the single-photon counting module instead of the spectrometer. UC intensity modulation faithfully follows that of applied voltage. We observed this overall behavior (fast reversible change of UC intensity) in 8 of the 28 NCs studied (Figure 4d). As the SHG peak did not vary at all in these cases during the voltage application and removal process (see the inset of Fig. 4b), we ruled out a change of focus of the microscope objective. Furthermore, the SHG peak stability also indicates that the applied electric field did not degrade the crystal lattice. Among these eight particles displaying a fast change in UC with constant SHG signal, seven of them reacted to only one polarity (among which nanoBTO:Yb,Er #1 and #2, Figure 4a,b) and a single one reacted to both polarities in an opposite manner. The responses of each of the eight nanocrystals to 10 V are summarized in Supporting Information Table S1.
@@
To determine whether the change in UC intensity in nanoBTO:Yb,Er is indeed attributed to the ferroelectric properties of the crystal host as reported for BTO:Yb,Er thin film 14, we conducted the following control experiments. First, we checked that the sole mechanical contact of the unbiased tip with the crystal does not induce a UC spectral change by piezoelectric effect (see Figure S4). Then, we considered erbium-doped nanocrystals of yttrium oxide (Y2O3:Er, with 2% doping). Y2O3 has a cubic crystallographic structure and therefore is devoid of piezoelectric and ferroelectric properties. Supporting Figure S5 shows that V applied to one Y2O3:Er3+ nanocrystal (kind gift of Diana Serrano, synthesis done by Shuping Liu and detailed in ref.34) did not induce any change of erbium up-conversion intensity. The absence of UC change under an applied voltage was confirmed for three other nanocrystals, evidencing that the ferroelectric BTO matrix is essential for the effect to happen.
@@
The 20 other NCs studied exhibited different behaviors. First, 11 particles did not show any UC or SHG variation whatever the polarity was (Figure S6a). Two displayed slow (few seconds) changes in UC with no change in SHG (Figure S6b); five displayed fast changes in UC and also in SHG (Figure S6c), while two showed a slow change in UC and SHG (Figure S6d). This diversity of behaviors is likely to reflect the one of the nanocrystal polarization texture combined with the one of the orientation of their crystallographic axis relative to the direction of applied electric field . These various configurations differ from the well defined one reported in Hao et al. 14, which involved BTO:Yb,Er doped thin film with oriented faces and parallel to . In this later case, applied parallel to (being the direction , indexed by 3) induces, by inverse piezoelectric effect, a strain , where is one of the piezoelectric coefficients of BTO. This strain is accompanied with ions relative displacement leading to variations of the crystal field around erbium and ytterbium ions, thereby changing the transition coefficients and subsequently UC intensity.
The case of BTO NCs is different to the one of BTO thin layer, as we recently evidenced that NCs do not have out-of-plane polarization in any of their facets 35. Indeed, we showed by simulations that the polarization along the polar axis rotates by 90∘ over about 10 nm, to become in-plane within the facets and prevent a strong electrostatic depolarizing field. We confirmed by piezoresponse force microscopy (PFM) that we detect only in-plane displacements and not any out-of-plane displacement. An electric field of amplitude applied to the facet by the PFM tip induces a shear strain displacement of amplitude , with pm/V being even larger than pm/V (that gives the strain amplitude in the case of BTO thin layer) according to Ref.36. This shear strain is also expected to induce crystal field deformations leading to UC intensity changes. Furthermore, our simulations showed also that the lateral piezoresponse amplitude is one order of magnitude larger in plane than in and planes 35. This result may explain why for a significant fraction of nanoBTO:Yb,Er (39%, “no change” in Figure 4d) we were not able to detect any PL changes, consistently with the fact that these particles were oriented with their axis perpendicular to . However, this interpretation would require to conduct correlated UC variation recordings and PFM measurements to be validated. The four other categories of particles for which PL change was observed (Figure 4d) encompass the case of aligned with (likely to be the situations yielding “large” PL change, as for the 8 particles displaying a fast change in UC with constant SHG), and perpendicular to but still yielding a detectable PL (case of particles with “large” size or having a polarization texture more favorable to larger electric field-induced strain). In a quarter of the cases, we also detected a reversible SHG intensity increase in the presence of the electric field (Figure S6c,d), which reflects crystal lattice deformation towards less symmetrical structure. Finally, we attribute slow UC response (4 particles, 14% of the cases, Figure S6b,d) to charge accumulation taking place when the conductive polymer glue does not fully ground the bottom part of the nanocrystal, but further research would be required to validate these two interpretations.

### v2

@@ Abstract
Understanding the dynamics of electrical signals within neuronal assemblies is crucial to unraveling complex brain function. Despite recent advances in employing optically active nanostructures in transmembrane potential sensing, there remains room for improvement in terms of response time and sensitivity. Here, we report the development of such a nanosensor capable of detecting electric fields with a submillisecond response time at the single particle level. We achieve this by using ferroelectric nanocrystals doped with rare earth ions producing upconversion (UC). When such a nanocrystal experiences a variation of surrounding electric potential, its surface charge density changes, inducing electric polarization modifications that vary, via converse piezoelectric effect, the crystal field around the ions. The latter variation is finally converted into UC spectral changes, enabling optical detection of electric potential. To develop such a sensor, we synthesized erbium and ytterbium-doped barium titanate crystals of size nm. We observed distinct changes in the UC spectrum when individual nanocrystals were subjected to an external field via a conductive AFM tip, with a response time of 100 s. Furthermore, our sensor exhibits a remarkable sensitivity of 4.8 kV/cm/, enabling time-resolved detection of fast changing electric field of amplitude comparable to that generated during a neuron action potential.
@@
Left: Electrical and optical scanning probe and fluorescence microscopy setup, to modulate and detect the up-conversion from single barium titanate nanocrystals doped with ytterbium and erbium ions (BaTiO3:Yb,Er). Right top: intensity of up-conversion (UC) under a voltage alternating between 0 and 10 V applied to a single nanocrystal, demonstrating the ability to sense such variations with UC at 100 µs temporal resolution. Right bottom: Spectral variations associated to applied voltage of 10 V.
@@ 1 Introduction
The generation and propagation of electrical signals in neurons and across neuronal assemblies, which occur at the nanometer and millisecond scales, remain poorly understood, despite the development of several theoretical models. For instance, the theory of core conductor cable and the Goldman-Hodgkin-Huxley-Katz models of ion motion and voltage propagation have provided theoretical foundations for neuronal physiology 1, 2. These theories account for the action potential, the few millisecond lasting voltage variation of 100 mV across the nm thick membrane (corresponding to 200 kV/cm peak voltage), which supports neuron-to-neuron communication. However, these models neglect electrodiffusion, that is, concentration changes associated with ionic currents. This assumption is true at millimeter scale like the squid giant axon but fails to describe accurately the processes taking place at the sub-micrometer size structure of synapses 3, 4. The Poisson-Nernst-Planck theory reveals that the geometry of synaptic compartments influences the distribution of local electric field generated during synaptic transmission, which subsequently changes the concentration of charged neurotransmitters undergoing electrodiffusion 5 and therefore has a functional impact. As the shape of the synapse changes during the execution of cognitive tasks 6, being able to measure the electric field on this scale would not only be an observable of such key plasticity but could provide an experimental test of the theory. There is thus a clear need to develop non-toxic, local methods to sense electric fields in biological media in order to refine current models of neuronal activity.
The detection of electric fields at the nanometer scale has initially been addressed using ultrahigh resolution nanopipette electrophysiology 7. Although very powerful, this technique can only study a single synapse at a time. On the contrary, voltage indicators (VI) were designed to stain the entire neuronal membrane and allow fluorescence recording of transmembrane voltage dynamics 8 throughout the cell with optical diffraction resolution. Enormous progress has been made in the development of brighter and faster VI, and in particular thanks to the development of hybrid chemigenetic systems that allow to combine optimally chemical compounds with genetically encoded proteins 9. However, subdiffraction imaging of a large number of synapses at high temporal resolution has not yet been reported. The development of a technology offering the necessary high spatiotemporal resolution (typically 1 ms, 200 nm) to reliably interrogate electrophysiological dynamics at multiple neuronal nanodomains would provide invaluable insights into open questions in neuronal electrophysiology.
@@
However, to maintain diffraction-limited imaging, an electric field must be sensed by a single or very few nanoparticles only. There are a few examples of such systems that are based on electric-field modulation of charge or Förster resonant energy transfer (FRET) between the nanoparticle and another entity. For example, Nag et al. used a QD-fullerene (C60) conjugate, with the C60 embedded in the cell membrane 12. At resting membrane potential the photoexcited QDs relax to the ground state partially by electron transfer to the C60. This transfer increases upon cell membrane depolarization, leading to a PL intensity relative quenching of a few % that provides a transmembrane potential variation readout both in cultured cell and in the cortex of live mouse. While chemical or electrical stimulation leads to QD PL modulation, the fastest response time was about 100 ms, two order of magnitude longer than the time response necessary to reproduce an action potential (AP) with fidelity. Another conjugate introduced by Liu et al. consists of an up-converting nanoparticle (UCNP) to which voltage sensing dyes were coupled, that serve on one hand to anchor the UCNP into cell membrane and are also FRET acceptors 13. The authors of this study report a relative change in UCNP luminescence intensity in the 0.4-6.1% range in cultured cells, depending on the stimulation conditions, and a time response of about 50 ms, still much longer than the duration of the AP. Hence, the spatial ( nm) and time (submillisecond) resolution required for membrane potential sensing remains to be addressed.
@@
Here, we report the development of nm-sized standalone optically active nano-crystals (NCs) capable of sensing, at the level of a single particle, a field of kV/cm slightly smaller than the peak amplitude field of an action potential, with a millisecond response time. We based our sensing principle on a specific UCNP made of a ferroelectric nanocrystal (NC) host doped with rare earth ions. When the NC is exposed to an external electric field, its lattice undergoes deformations due to a converse piezoelectric effect, which leads to spectral changes of the rare-earth ion up-conversion emission. The effectiveness of this principle was established by Hao et al. who demonstrated, by a varying electric field of 125 kV/cm amplitude, the modulation of the upconverted (UC) light from an ytterbium (Yb3+) and erbium (Er3+)-doped epitaxial barium titanate (BTO) thin film 14, at room temperature (in BTO ferroelectric phase). The authors demonstrated an UC modulation at a frequency of 0.1 Hz, far below the maximum frequency of kHz limited by the UC process response time of s in erbium-doped BTO 15. Thus, here we develop BTO:Yb,Er individual nanocrystals of size nm, which we expose to a fast varying electric field with the conductive tip of an atomic force microscope (AFM), coupled to a fluorescence microscope. We demonstrate that 47% of the nanocrystals exhibit a fast variation in UC intensity with a response time of s, limited by the duration of the UC optical cycle. Furthermore, BTO nanocrystals are ideal to develop sensors that could be rapidly tested in biological contexts due to their pre-existing use in drug delivery or tissue engineering 16, but also for their intrinsic optical two-photon excitation response17 (resulting from the crystal second-order nonlinear response) in cultured cells 18, 19, 20 and larvae of small organisms 21, 22, without or with grafted functional groups 22, 19, 20.
@@ 2.1 Synthesis and characterization of erbium and ytterbium-doped BTO nanocrystals
Erbium and ytterbium-doped BTO nanocrystals (nanoBTO:Yb,Er) were synthesized by a combination of coprecipitation and hydrothermal methods. Ytterbium was used to sensitize erbium up-conversion. By adapting the Ba/Ti stoichiometry, we aimed at incorporating Er3+ and Yb3+ ions in A and B sites of the ABO3 barium titanate perovskite structure respectively, because the ytterbium ions have the closest radius to the one of titanium (B site) in the six coordinated state 23. We targeted the stoichiometry , ensuring charge compensation, as described in Materials and Methods. The choice of an erbium concentration of 1% was made based on Zhang et al.24, who showed that higher concentrations enhance cross-relaxation, leading to lower UC.
@@
One may wonder where Er3+ and Yb3+ ions insert themselves in the BTO nanocrystal. Previous works 29, 30 indicate that Yb3+ ions predominantly substitute Ti4+ ions, whereas Er3+ can occupy both sites. For charge compensation reasons, it is thus very likely that Er3+ substitutes Ba2+, but we cannot exclude that they also fill B sites 31, which would lead to the appearance of oxygen vacancies. To further investigate effective substitution sites and the presence of such defects, other studies, such as UC lifetime 15 and conductive measurements 32 would be necessary but are beyond the scope of this article.
Increasing the effective content in Yb3+ may lead to brighter NC. However, despite our efforts which consisted in varying multiple parameters during the synthesis (Ba/Ti ratio, known to influence Er3+ site of incorporation 31; erbium precursor compound; and ions concentrations between 0.25 and 2% for Er3+ and between 2.5% and 13% for Yb3+), we were not able to introduce a larger proportion of ytterbium. The sample selected for the present study to test the electric field sensing is the brightest we achieved among the different synthesis we made. We noticed in high-resolution STEM of samples dimmer than the one selected for this article, the presence of polycrystalline ytterbium oxide nanoparticles surrounding the nanoBTO, which is consistent with a segregation of ytterbium outside the BTO host.
The characterization studies validated that the synthesized BTO nanocrystals are doped with ytterbium and erbium, and keep the tetragonal structure of undoped nanocrystals. Hence, nanoBTO:Yb,Er are very likely to maintain the ferroelectric and piezoelectric properties on which we based the electric field sensing mechanism. We then investigated the upconversion photoluminescence (PL) of individual nanoBTO:Yb,Er, and the modulation of its intensity by an external electric field.
@@ 2.2 Up-conversion of single nanoBTO:Yb,Er and its modification by an electric potential
Note that 4f-4f electric dipole transitions of lanthanide ions are expected to be forbidden by quantum mechanics selection rules in a perfectly centrosymmetric crystalline environment, for which the odd orders of the crystal field vanish. The mere detection of erbium UC, even in absence of an external electric field, is the signature of lower local symmetry around erbium ions. In fact, the splitting of the Miller indexes (002) and (200) Bragg peaks near in the inset of Fig. 1a already indicates the presence of the polar (hence non-centrosymmetric) phase of BTO, enabling the 4f-4f electric dipole transitions of lanthanide ions. Note also that Zou et al. 33 reported erbium UC even in the high-temperature centrosymmetric phase of BTO, which indicates that the local environment felt by the RE ions is likely even less symmetric than the symmetry of the overall matrix, which is consistent with the eight-site model developed in BTO 34 in which local symmetry breakings exist at the unit cell level but vanish on a macroscopic scale.
@@
We then investigated the changes of UC spectrum of a single nanoBTO:Yb,Er (nanoBTO:Yb,Er #1) when we apply a static voltage of V across the particle. For the average particle size of 158 nm, this voltage corresponds to an electric field of kV/cm, more than twice smaller than the electric field breakdown of 1.5-3 MV/cm reported for thin BaTiO3 in the literature 35. Figure 4a shows UC spectral changes of nanoBTO:Yb,Er #1 under static voltage exposure of +10 V, manifested by an overall increase in the intensity of all peaks. This variation was reversible when the voltage was returned to 0 V, reverting to its initial value, and no change was detected with the opposite polarity of -10 V. For another NC (nanoBTO:Yb,Er #2), we also detected a reversible UC emission increase under an opposite applied voltage of -10 V (Figure 4b), also with no change for the opposite polarity (of +10 V). The intensities of the green and red UC emission bands both increase with the applied voltage, regardless of its polarity. This behavior differs from the one reported in BTO thin layer by Hao et al.14, where they observed an increase in the intensity of the green bands but none of the red one. Hao et al. proposed an interpretation based on the transition probability predicted by Judd–Ofelt theory, where the dominant contribution for the green emission transitions involves the intensity parameter that increases when the symmetry of the site occupied by Er3+ is reduced, which is what happens when the crystal is exposed to the external electric field. In contrary, the red emission transition does not involve this symmetry sensitive term and therefore should be almost insensitive to the electric field. Our observation of a variation of intensity of the red emission transition under an external electric field indicates that in our sample, there is a more complex coupling of the red transition to the site symmetry changes. We hypothesize that the change of intensity of the red line could be due to an increase of population of the 4F9/2 upper level of the red transition resulting from a larger cross-relaxation effect. The latter depends on the distance between ions, which indeed is modulated by the lattice deformation resulting from the converse piezoelectric effect, under the applied electric field. Indeed we evidenced that in our nanoBTO:Yb,Er sample the red transition is governed by the cross-relaxation, because we observed that its intensity evolves linearly with the excitation laser intensity (see Figure S3), consistently with two red photons emitted for two infrared excitation photons absorbed, with one of them experiencing a cross relaxation (see Figure 3d).
@@
As for NC like nanoBTO:Yb,Er #1 and #2 the intensities of both the green and red bands increase similarly under applied electric field, we considered the total UC intensity integrated over the whole spectrum. Figure 4c displays this total UC intensity change under a voltage square modulation for nanoBTO:Yb,Er #1, if whole the UC light is sent to the single-photon counting module instead of the spectrometer. UC intensity modulation faithfully follows that of applied voltage. We observed this overall behavior (fast reversible change of UC intensity) in 8 of the 28 NCs studied (Figure 4d). As the SHG peak did not vary at all in these cases during the voltage application and removal process (see the inset of Fig. 4b), we ruled out a change of focus of the microscope objective. Furthermore, the SHG peak stability also indicates that the applied electric field did not degrade the crystal lattice. Among these eight particles displaying a fast change in UC with constant SHG signal, seven of them reacted to only one polarity (among which nanoBTO:Yb,Er #1 and #2, Figure 4a,b) and a single one reacted to both polarities in an opposite manner. The responses of each of the eight nanocrystals to 10 V are summarized in Supporting Information Table S1.
@@
To determine whether the change in UC intensity in nanoBTO:Yb,Er is indeed attributed to the ferroelectric properties of the crystal host as reported for BTO:Yb,Er thin film 14, we conducted the following control experiments. First, we checked that the sole mechanical contact of the unbiased tip with the crystal does not induce a UC spectral change by piezoelectric effect (see Figure S4). Then, we considered erbium-doped nanocrystals of yttrium oxide (Y2O3:Er, with 2% doping). Y2O3 has a cubic crystallographic structure and therefore is devoid of piezoelectric and ferroelectric properties. Supporting Figure S5 shows that V applied to one Y2O3:Er3+ nanocrystal (kind gift of Diana Serrano, synthesis done by Shuping Liu and detailed in Ref.36) did not induce any change of erbium up-conversion intensity. The absence of UC change under an applied voltage was confirmed for three other nanocrystals, evidencing that the ferroelectric BTO matrix is essential for the effect to happen. Finally, as one may wonder if the co-doping with Yb3+ may play a key role in the transduction mechanism, we also tested NaYF4 nanoparticles doped with 2% erbium and 18% ytterbium (kind gift of Thierry Gacoin, synthesis by Godefroy Leménager detailed in Ref.37). We used nanocrystals in the -NaYF4 phase which structure belongs to the P63/m space group. The latter has polar properties, and therefore a NaYF4 NC may also deform when exposed to an electric field. However, we did not observe any modulation of its UC under a variable applied voltage, as shown in Supporting Figure S6. Hence, this additional experiment demonstrates the superior properties of ferroelectric BaTiO3 NCs over another polar material such as NaYF4:Yb,Er nanocrystals.
@@
The 20 other NCs studied exhibited different behaviors. First, 11 particles did not show any UC or SHG variation whatever the polarity was (Figure S7a). In this case, bad physical contact between the tip and the NC top facet was ruled out because we engaged and withdrew the tip several times and obtained the same negative result. Two NC displayed slow (few seconds) changes in UC with no change in SHG (Figure S7b); five displayed fast changes in UC and also in SHG (Figure S7c), while two showed a slow change in UC and SHG (Figure S7d). Note that we did not observe any correlation between the nanocrystals size and UC response: nanoBTO:Yb,Er#1 to #3, that displayed UC changes upon applied voltage variation, have measured height of 153, 111 and 130 nm respectively, while nanoBTO:Yb,Er of Figure S7a, which did not manifest UC changes, has a height of 133 nm, similar to the one of reacting NC.
The diversity of behaviors could rather have its origin in at least two phenomena: electrical current leakage through the nanocrystal, preventing the electric field to be applied, and/or the complex nanocrystal polarization texture that renders its crystal deformation under dependent on the orientation of the field relative to the NC crystallographic axis as we recently evidenced in Ref. 38. Current leakage has been reported in rare earth-doped ferroelectric thin films and was attributed to mobile oxygen vacancies appearing to ensure charge compensation. More precisely, it has been shown in (Ba,Sr)TiO3 thin films, that when doping with a RE3+ ion only on the A2+ or on B4+ site, the resulting charge imbalance generates oxygen vacancies responsible for a large increase in leakage 32. We hypothesize that the lack of electric field modulation of the UC may be due entirely or partly to a similar mechanism where Er3+ and/or Yb3+ mainly dope either the A2+ or B4+ site, but not both at the same time. Further experiments involving conductive-AFM and nanoscale analytical techniques (Electron Energy Loss Spectroscopy, EDX or Atom Probe Tomography 39, 40), outside of the scope of this article, are needed to test this hypothesis.
@@
As for the other source of variable response related to the polarization texture, we have to take into account the orientation of relative to the NC crystallographic axis. Indeed, our geometrical configuration differs from the well defined one reported in Hao et al. 14, which involved BTO:Yb,Er doped thin film with oriented faces and parallel to . In this later case, applied parallel to (being the direction , indexed by 3) induces, by inverse piezoelectric effect, a strain , where is one of the piezoelectric coefficients of BTO. This strain is accompanied with ions relative displacement leading to variations of the crystal field around erbium and ytterbium ions, thereby changing the transition coefficients and subsequently UC intensity.
In the case of BTO NCs we recently showed by simulations and piezoresponse force microscopy that the polarization along the polar axis rotates by 90∘ over about 10 nm, to become in-plane within the facets, preventing a strong electrostatic depolarizing field. An electric field of amplitude applied to the facet by the AFM tip induces a shear strain displacement of amplitude , with pm/V being even larger than pm/V, according to Ref.41. This shear strain is also expected to induce crystal field deformations leading to UC intensity changes. Furthermore, our simulations showed also that the lateral piezoresponse amplitude is one order of magnitude larger for perpendicular to plane than to or planes 38. This dependence of the deformation amplitude to the electric field orientation relative to the axis, may contribute to the diversity of UC variation observed in the present work.
Polarization-resolved second-harmonic generation radiative pattern measurement is conventionally used to determine the crystalline axis orientation of a monocrystal, even nanometric size ones 42. However, this method does not work with BTO nanocrystals due to their complex polarization texture 38. Hence we are not able to determine the axis orientation of the individual NC probed for their electric-field sensing ability, which prevents us to check the influence of orientation relative to on the signal intensity.